OpenStudy (anonymous):

Find all solutions in the interval [0,2(pi)] sin2x=cosx

OpenStudy (anonymous):

You can rewrite sin2x using the double angle formula to get: cosx = 2*sinx*cosx... subtract cosx on both sides 2sinxcosx - cosx = 0... factor cosx(2sinx - 1) = 0 Set each factor equal to zero to get two sets of roots: cosx = 0 ==> x = {π/2,3π/2} 2sinx - 1 = 0 ==> sinx = 1/2 ==> x = {π/6, 5π/6}

OpenStudy (anonymous):

\[\sin(2x)=2\sin(x)\cos(x)\] so put \[2\sin(x)\cos(x)=\cos(x)\] \[2\sin(x)\cos(x)-\cos(x)=0\] \[\cos(x)(2\sin(x)-1)=0\] \[\cos(x)=0\] or \[2\sin(x)-1=0\]

OpenStudy (anonymous):

damn i am slow

OpenStudy (anonymous):

slowpoke

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

lol. but i am using latex!

OpenStudy (anonymous):

stop making excuses